OpenStudy (anonymous):

What is the equation of the line that passes through the points (1,2/3) and (-1,-1) ? A.y=-6x+5 B.y=5/6x-1/6 C.y=2/3x+1/5 D.y=6/5x+1/5

hartnn (hartnn):

can you find the slope of line first ?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

alright thank you :)

hartnn (hartnn):

after you get the slope,m you use \(y-y_1=m(x-x_1)\) where x1,y1 is any one of the 2 points...

OpenStudy (anonymous):

alright i understand it a little better now :)